### v2

# HueManity: Probing Fine-Grained Visual Perception in MLLMs

Abstract
Multimodal Large Language Models (MLLMs) excel at high-level visual reasoning, but their performance on nuanced perceptual tasks remains surprisingly limited. We present HueManity, a benchmark designed to assess visual perception in MLLMs. The dataset comprises 83,850 images featuring two-character alphanumeric strings embedded in Ishihara test style dot patterns, challenging models on precise pattern recognition. Our evaluation of nine state-of-the-art MLLMs on HueManity demonstrates a significant performance deficit compared to human and traditional computer vision baselines. The best-performing MLLM achieved a 33.6% accuracy on the numeric ‘easy’ task and a striking 3% on the alphanumeric ‘hard’ task. In contrast, human participants achieved near-perfect scores (100% and 95.6%) [^1] Human evaluations utilized 100-image representative subsets, sampled from the model evaluation sets for each respective task., and a fine-tuned ResNet50 model reached accuracies of 96.5% and 94.5%. These results highlight a critical gap in the visual capabilities of current MLLMs. Our analysis further explores potential architectural and training-paradigm factors contributing to this perceptual gap in MLLMs. We open-source HueManity dataset and code to foster further research in improving perceptual robustness of MLLMs.
Code: https://github.com/rynaa/huemanity Dataset: https://huggingface.co/datasets/Jayant-Sravan/HueManity
HueManity: Probing Fine-Grained Visual Perception in MLLMs
Rynaa Groverthanks: Equal Contribution Google rynaa@google.com                                   Jayant Sravan Tamarapalli11footnotemark: 1 Google jayantsravan@google.com
Sahiti Yerramilli11footnotemark: 1 Google sahitiy@google.com                                   Nilay Pande11footnotemark: 1 Waymo nilayp@waymo.com
Figure 1: HueManity — A new benchmark for MLLM fine-grained visual perception. The pipeline creates a character mask from alphanumeric characters and renders it as an Ishihara-style pattern. While models achieve high accuracy on clear masks, they struggle significantly on the challenging pattern images.
1 Introduction
The trajectory of Multimodal Large Language Models (MLLMs) Team et al. (2023); Achiam et al. (2023); Bai et al. (2023a); Li et al. (2023b); Gong et al. (2023); Liu et al. (2024a, 2023); Anthropic (2025) in recent years has been marked by impressive advancements, demonstrating sophisticated capabilities in bridging visual and textual information. Their capabilities extend well beyond simple image labeling Russakovsky et al. (2015); Deng (2012), enabling complex tasks like generating detailed image descriptions Dong et al. (2024); Fu et al. (2024a), answering intricate visual questions requiring inference about relationships and activities Weng et al. (2025); Chen et al. (2025); Kuang et al. (2024), and participating in nuanced dialogue about visual content Cao et al. (2024). A key factor in their success is pre-training on vast web-scale image-text datasets, which facilitates learning powerful representations that capture high-level semantic links between visual features and language Jia et al. (2021); Radford et al. (2021); Schuhmann et al. (2022); Alayrac et al. (2022); Qi et al. (2020); Zhai et al. (2022); Pham et al. (2023). Consequently, MLLMs perform strongly on tasks benefiting from this conceptual understanding, including recognizing common objects, interpreting general scene structure, and relating them to text.
However, the predominant evaluation paradigms for MLLMs have centered on these conceptual capabilities, largely overlooking their fine-grained perceptual acuity Bai et al. (2023b); Li et al. (2024, 2023a); Xu et al. (2024); Yin et al. (2023); Liu et al. (2023). This leaves a critical gap in understanding their performance on tasks demanding precise visual discernment – such as intricate pattern recognition, subtle feature differentiation, and robust segregation within visually cluttered background. Unlike tasks solvable through broad semantic association, these perceptual challenges require a more fundamental visual processing capacity, akin to the human ability to resolve intricate details. This paper introduces a benchmark specifically designed to probe this nuanced dimension of MLLM performance. Our methodology draws inspiration from the principles of Ishihara plates Clark (1924), a technique traditionally employed in human ophthalmology to assess color vision by embedding figures (like numbers or paths) within fields of multicolored, varied-size dots. It is crucial to clarify that HueManity does not aim to diagnose ‘color blindness’ in MLLMs. Instead, our Ishihara-style stimuli, created using controlled generation techniques, rigorously test an MLLM’s fundamental ability to identify embedded alphanumeric characters by their subtle color and luminance contrasts within visually cluttered dot patterns.
Success on the HueManity benchmark serves as a crucial indicator of an MLLM’s potential for robust visual understanding in complex, real-world scenarios. Unlike often curated benchmark datasets, real-world visual environments are frequently characterized by clutter, partial occlusions, variable lighting, and unconventional information presentation. The ability of an MLLM to reliably parse characters in our Ishihara-style plates is intended to assess its resilience to visual clutter and its pattern recognition capabilities — foundational skills often linked to dependable performance in challenging visual settings. Furthermore, mastering this level of fine-grained perception is foundational for more intricate forms of visual reasoning. An MLLM’s struggles with basic pattern recognition under challenging conditions could imply difficulties in performing higher-order tasks that depend on the accurate interpretation of subtle details. Thus, HueManity serves not merely as a test of pattern recognition, but as a probe into architectural or training limitations that hinder MLLMs from achieving comprehensive, detailed visual intelligence.
To address this identified gap and facilitate further research in this domain, this paper makes the following specific contributions:
  1. 1.
    We introduce HueManity, a new large-scale benchmark (83,850 images) featuring Ishihara-inspired alphanumeric stimuli. The benchmark utilizes a principled design with 25 carefully curated color pairs, selected using CIEDE2000 ($\Delta E_{2000}$) metrics and manual verification, ensuring both systematic challenge and fairness for human comparison.
  2. 2.
    We conducted a comprehensive evaluation of nine state-of-the-art MLLMs on HueManity’s easy and hard sets. Findings reveal a significant performance gap compared to strong human and fine-tuned ResNet50 baselines. Baseline success suggests MLLM limitations are due to their current architectures or training for fine-grained perception, rather than the task being intractable.
  3. 3.
    We release open-source code for generating challenging Ishihara-style perceptual stimuli (alphanumeric characters embedded in dot patterns), enabling reproducible research and community-driven extensions.
2 Related Works
2.1 Multimodal Models
With the remarkable advancements of Large Language Models (LLMs), recent research has extended their capabilities to multimodal domains by integrating visual information, giving rise to Multimodal Large Language Models (MLLMs) Team et al. (2023); Achiam et al. (2023); Bai et al. (2023a); Li et al. (2023b); Gong et al. (2023); Liu et al. (2024a, 2023). These models typically align visual features from pre-trained image encoders with LLMs via modality adaptation layers. Early works like BLIP-2 Li et al. (2023b) pioneered this architecture by first pre-training on image-text datasets and fine-tuning on task-specific benchmarks such as Visual Question Answering (VQA). Subsequent models like LLaVA Liu et al. (2023) advanced this approach by leveraging synthetic instruction-following data in VQA formats, significantly improving instruction tuning performance. More recent efforts have expanded into video understanding and even image generation, showcasing the versatility of MLLMs across modalities. However, much of their celebrated success in visual tasks often appears intertwined with, and perhaps reliant upon, their powerful language capabilities to interpret, reason about, and generate text from visual content. This inherent strength in linguistic processing may have led to less emphasis on developing or scrutinizing their more fundamental, non-linguistic visual perception skills. HueManity directly addresses this gap in scrutinizing MLLM visual perception. It is a benchmark specifically designed to isolate and rigorously evaluate these core abilities.
2.2 MLLM Evaluation
Multimodal Large Language Models (MLLMs) are highly capable of global image understanding and reasoning but consistently struggle with fine-grained visual tasks such as precise recognition and localization Huang and Zhang (2024); Li et al. (2024). Several benchmarks have been proposed to address these challenges. TouchStone Bai et al. (2023b) offers 908 manually annotated visual dialog questions across five abilities and 27 sub-tasks, while LLaVA-Bench Liu et al. (2023) includes 24 images with 60 curated questions covering diverse content like scenes, memes, and sketches. To automate evaluation, LLaVA-Bench Liu et al. (2023), LAMM Yin et al. (2023), and TouchStone Bai et al. (2023b) rely on GPT-based models to judge relevance and accuracy, but this introduces inherent reliability issues and cost inefficiencies. Similarly, LVLM-eHub Xu et al. (2024) aggregates multiple vision benchmarks but still depends on human annotators to compare model outputs, making it both subjective and expensive. More structured efforts like MME and MMBench Liu et al. (2024b) attempt to provide objective evaluations with multiple-choice questions across a wide range of ability dimensions. MME pioneered multimodal Yes/No questions for perception and reasoning, with MM-Vet Yu et al. (2023) and MMBench Liu et al. (2024b) extending coverage to sub-tasks like OCR, math, and recognition, yet they all heavily rely on existing VQA datasets or GPT-generated questions. SEED-Bench Li et al. (2024, 2023a) scales up to 24,000 human-annotated multiple-choice questions covering varied input-output modalities, but its challenges remain simple, with most open-source models reaching 30–60% accuracy at the easiest level. Blink Fu et al. (2024b) attempts a more holistic evaluation across 14 perception tasks with 3,900 questions and 7,300 images, but does not assess models on combinations of tasks, and remains less challenging, as evidenced by higher model accuracies.
HueManity uniquely evaluates Multimodal Large Language Models (MLLMs) on the foundational visual skill of discerning patterns from subtle cues in cluttered environments, employing alphanumeric characters within Ishihara-style dot patterns with controlled color contrasts as a robust proxy. This stimulus design, combined with procedural generation and exact-match evaluation, establishes a scalable, objective, and reliable methodology distinct from more subjective or resource-intensive techniques.
3 Data Creation
The HueManity dataset comprises 83,850 images, each pairing a two-character alphanumeric string with its ground truth label and full generation parameters. These strings are formed from lowercase letters (a-z), uppercase letters (A-Z), and digits (0-9). To enhance clarity and prevent evaluation errors, we excluded visually ambiguous characters (‘l’, ‘I’, ‘J’, ‘O’) and combinations commencing with ‘0’ (e.g., ‘01’, to avoid ‘1’ vs. ‘01’ prediction conflicts). The entire dataset was produced by rendering all valid two-character combinations using 25 meticulously curated color pairs, which then formed the basis for our evaluation sets.
The full HueManity dataset, encompassing all 83,850 generated images, serves as a comprehensive resource for evaluating MLLM perception. However, for the specific MLLM evaluations reported in this paper, particularly considering practical constraints such as time and API access costs, we defined two distinct tasks using carefully sampled subsets. Each subset comprises 1,000 images, randomly selected to be representative for these focused evaluations:
  1. 1.
    Number Recognition Set (Used for the ‘Easier’ Task): This subset of 1,000 images is used to evaluate models on the Number Identification task. It exclusively contains images displaying two-character strings composed solely of numeric digits (e.g., ‘17’, ‘83’, ‘05’).
  2. 2.
    Text Recognition Set (Used for the ‘Harder’ Task): This subset of 1,000 images, drawn from the broader pool of alphanumeric combinations, is employed for the Text Identification task. It features diverse two-character alphanumeric strings, including various combinations of letters and digits (e.g., ‘A7’, ‘b9’, ‘XG’).
3.1 Text Mask Generator
The initial stage in our data pipeline is the creation of a 900x900 pixel resolution binary text mask for each two-character string. Using Pygame, we render these strings in white against a black background. To ensure both clear readability and adequate character width suitable for the subsequent dot-based rendering, all masks utilize the DejaVu Sans font (size 550, bold and italic styles), as exemplified in Figure 2(a).
(a) Binary Text Mask
(b) Ishihara pattern
Figure 2: Image generation process for the alphanumeric string “Y8”. (a) The binary text mask. (b) The corresponding Ishihara-style dot pattern.
3.2 Ishihara-Style Pattern Generation
Our pattern generator, adapted from an open-source Pygame project [^2] https://github.com/hakrackete/Ishihara-color-plate-generator, iteratively populates the image with non-overlapping circles. Over 30,000 iterations, the generator performs the following steps for each potential circle:
  1. 1.
    Placement: Randomly sample (x,y) coordinates and compute the maximum non-colliding radius (ranging from 4 to 15 pixels).
  2. 2.
    Initial Color: Assign initial color (foreground/background) based on whether the center falls within the character mask (Fig. 2(a)).
  3. 3.
    Color Transformation: Apply three randomized transformations to the initial color: gradient shift towards the other pair color, RGB color shift (range $[-30,+30]$), and RGB lightness scaling (factor 0.66-1.5).
  4. 4.
    Rendering: Render the transformed circle at the computed position.
This iterative process ultimately yields the final, dense Ishihara-style pattern (as shown in Figure 2(b)).
3.3 Color Pairs Selection
The 25 distinct foreground-background color pairs used in the HueManity stimuli were meticulously selected through a multi-stage process to ensure a balance between perceptual challenge and unambiguous human legibility. This procedure involved quantitative CIEDE2000 Luo et al. (2001) analysis and extensive manual verification, with the full methodology detailed in Appendix B.
4 Experiments
We evaluated a diverse set of nine state-of-the-art Multimodal Large Language Models (MLLMs), encompassing prominent commercial APIs and publicly available open-source models. Model inference was orchestrated using Promptfoo [^3] https://www.promptfoo.dev, a platform facilitating reproducible benchmarking through flexible prompt definition and API integration. Open-source models were hosted locally via Ollama [^4] https://ollama.com and inferred on a single NVIDIA A100 GPU.
Evaluation focused on two primary tasks: numerical recognition (digits only) and alphanumeric recognition (digits and letters). For each of these tasks, a subset of 1,000 points from HueManity: two-character strings and their corresponding generated images were used. Models were evaluated on two types of visual stimuli for each datapoint:
  • HueManity Plates: The 1,000 randomly sampled Ishihara-style dot pattern images from the HueManity benchmark.
  • Text Masks: The corresponding 1,000 binary text mask images (i.e., white text on a black background, as exemplified in Figure 2(a)). This additional evaluation on text masks was performed to establish a baseline for each model’s ability to recognize the characters in a clear, unobstructed format, helping to differentiate fundamental OCR capabilities from performance on the perceptually challenging dot patterns.
All images (both HueManity plates and text masks) were provided to the models encoded in Base64 format. The following task-specific prompts were used for both stimulus types:
  • Number Recognition Prompt: "What is the number in this image? Strictly stick to the format: Answer: [number in the image]"
  • Text Recognition Prompt: "What is the exact text in this image? It has only alpha-numeric characters excluding small L, capital O, capital I, and capital J to avoid ambiguity. Strictly stick to the format: Answer: [exact text in the image]"
4.1 Human Performance Evaluations
To establish a human performance baseline, evaluations involved three adult volunteers, all with self-reported normal color vision. This group size was considered adequate because identifying characters in HueManity plates is a rudimentary perceptual task for humans with normal vision, one that requires no specialized skills and typically exhibits low inter-volunteer performance variance. For each of the two tasks (numerical and alphanumeric recognition), these volunteers were tested on a representative subset of 100 images, randomly sampled from the corresponding 1,000-image evaluation sets used for MLLMs.
Volunteers viewed the images at their original 900x900 pixel resolution within a Google Sheets document. Using the identical prompts provided to the MLLMs, they were asked to identify the embedded characters, entering their responses directly into the document. This methodology ensured a direct comparison of human and MLLM performance under equivalent conditions.
4.2 Traditional Computer Vision Baseline (ResNet50)
As a representative traditional computer vision baseline, we trained and evaluated a ResNet50 model. We utilized a ResNet50 pre-trained on ImageNet, obtained from the PyTorch vision library.
The standard classification layer of the ResNet50 was replaced with two independent classification heads. Each head was designed to predict one character, treating the task as two independent character recognition problems. For the purpose of fine-tuning this model, we utilized 2,000 images randomly sampled from the broader HueManity dataset, ensuring these were distinct from the final evaluation subsets. Training was conducted for 30 epochs using the Adam optimizer with a learning rate of $31e-$ 3. The loss function for training was the sum of the cross-entropy losses calculated independently for each of the two classification heads. The trained model was evaluated on the same 1,000-image subsets used for the MLLM evaluations.
5 Results and Analysis
5.1 Human Performance: A Near-Perfect Baseline
Human volunteers established a crucial performance baseline, demonstrating exceptionally high accuracy and efficiency on the HueManity benchmark (detailed in Table 3). On the numerical recognition task, volunteers achieved a perfect 100% average accuracy, indicating that identifying two-digit strings within the dot patterns was straightforward. The more complex alphanumeric task also yielded a very strong average accuracy of 95.6%.
The minor errors in the alphanumeric task were insightful: annotator feedback consistently attributed confusion not to an inability to perceive the characters, but to difficulty differentiating visually similar character forms (e.g., ‘s’, ‘c’, ‘w’ upper and lower case variants). This distinction is important, as it underscores that the primary perceptual challenge of extracting patterns from the dots was readily overcome by the human visual system. Furthermore, volunteers processed these images with remarkable efficiency, typically requiring less than one second per image to identify the characters.
Near-perfect and rapid human scores offer crucial insights into MLLM performance on this task for several key reasons:
  • Confirms Task Solvability and Stimulus Clarity: High human accuracy confirms the stimuli provide unambiguous visual information for character identification, showing the tasks are well-defined and solvable by proficient visual systems.
  • Establishes a Performance Ceiling: Human performance sets a clear accuracy ceiling, highlighting the target for machine perception systems on this (for humans) perceptually simple task.
  • Highlights the Nature of MLLM Deficiencies: Significant MLLM deviations from human scores indicate machine-specific limitations, not task impossibility or unclear stimuli. This contrast helps determine if MLLM struggles lie in initial perceptual grouping or later recognition.
  • Contextualizes Machine Errors: Human ease and minor confusions (e.g., similar glyphs) provide valuable context for analyzing MLLM error types and severity, pinpointing where machine perception significantly diverges from human visual understanding.
In essence, the human baseline shows HueManity tests foundational visual skills that humans execute efficiently, despite MLLM challenges. This context is vital for appreciating the scale and nature of MLLM performance gaps.
5.2 ResNet50 Baseline: Demonstrating Task Learnability
To further contextualize the performance of Multimodal Large Language Models (MLLMs), we established a strong baseline using a fine-tuned ResNet50 He et al. (2015), a well-established traditional computer vision architecture. As detailed in Table 1 and Table 2, the ResNet50 model achieved high accuracy, scoring 96.5% on the numerical recognition task and 94.5% on the alphanumeric task.
This robust performance from a standard convolutional architecture, fine-tuned on a relatively small set of 2,000 images from the HueManity dataset (as described in Section 4.2), is highly informative. It clearly demonstrates that the task of identifying characters within the HueManity dot patterns is fundamentally learnable by established computer vision techniques.
Furthermore, achieving near-human accuracy with this setup suggests that the perceptual cues embedded in the stimuli are sufficiently rich and consistent for a focused model to learn effective recognition. This implies that the task, while designed to challenge nuanced perception, is not inherently intractable for AI and can be considered relatively easy for appropriately adapted vision models. The ResNet50 baseline’s success, therefore, pinpoints the challenges as MLLM-specific. This suggests that MLLM difficulties likely arise not from an inherently unsolvable visual task, but from how these larger, more general models currently process or prioritize fine-grained perceptual information.
Table 1: Accuracy on the number recognition task for human evaluators, ResNet50, and various MLLMs on both text masks and patterned HueManity images.
Mask Patterned Image
Humans (average) 100% 100%
ResNet50 - 96.5%
API-based models
GPT-4.1 mini 100% 19.0%
GPT-4.1 100% 33.6%
Claude 3.7 Sonnet 100% 0.4%
Open-source models
LLaVA-v1.6-7B 87.7% 3.3%
LLaVA-v1.6-13B 87.2% 8.1%
LLaVA-v1.6-34B 96.6% 7.8%
Mistral-small3.1-24b 100% 0.1%
Qwen VL Max 100% 0.2%
Pixtral 100% 1%
Table 2: Accuracy on the alphanumeric recognition task, comparing human performance with ResNet50 and MLLMs across text masks and patterned HueManity images.
Mask Patterned Image
Humans (average) 100% 95.6%
ResNet50 - 94.5%
API-based models
GPT-4.1 mini 72.4% 0.6%
GPT-4.1 80% 3.0%
Claude 3.7 Sonnet 82.2% 0%
Open-source models
LLaVA-v1.6-7B 15% 0%
LLaVA-v1.6-13B 31.8% 0.1%
LLaVA-v1.6-34B 27.1% 0%
Mistral-small3.1-24b 58.7% 0%
Qwen VL Max 83.5% 0%
Pixtral 65.8% 1.8%
Table 3: Human evaluation accuracies on 100-image subsets for both number and alphanumeric pattern recognition tasks.
Number Pattern Alphanumeric Pattern
Evaluator 1 100% 97%
Evaluator 2 100% 95%
Evaluator 3 100% 95%
Average 100% 95.6%
5.3 The Perceptual Gap: MLLM Performance vs. Baselines on HueManity
The performance of the nine evaluated Multimodal Large Language Models (MLLMs) on the HueManity benchmark stands in stark contrast to the near-perfect accuracies achieved by both human volunteers and the fine-tuned ResNet50 baseline He et al. (2015) (detailed in Table 1 and Table 2). Across both the numerical and alphanumeric tasks, all evaluated MLLMs exhibited a profound and consistent inability to match these baselines. For instance, even the best-performing MLLM achieved only 33.6% on the easier numeric task and a mere 3% on the harder alphanumeric task—figures that dramatically underscore their struggle.
This significant and consistent underperformance, on tasks demonstrably solvable by human vision and traditional AI methodologies, signals a critical gap in the fine-grained visual perception capabilities of current MLLMs. Such results necessitate a deeper exploration into the potential underlying reasons for these models’ difficulties.
Several characteristics inherent to the current design and training paradigms of many Multimodal Large Language Models (MLLMs) may contribute to their observed difficulties on tasks demanding nuanced visual perception. Firstly, the visual processing pipeline itself could be a factor. Vision encoders in MLLMs, though often powerful and pre-trained on diverse datasets (e.g., ViT variants Liu et al. (2023, 2024a); Agrawal et al. (2024); Bai et al. (2023a, 2025), are typically optimized for capturing semantic information and global scene context. This optimization focus might lead to the unintentional down-weighting or loss of extremely fine-grained local details—like subtle color shifts defining patterns within a dense, similarly featured background—that are crucial when strong semantic cues are absent. Furthermore, the interface used to connect these vision encoders to the language model—often a projection layer or a small MLP Liu et al. (2023, 2024a); Agrawal et al. (2024); Bai et al. (2023a, 2025)—necessarily transforms and condenses visual information. This critical step, while enabling multimodal fusion, might inadvertently act as an information bottleneck, potentially abstracting or losing the precise, high-resolution feature distinctions necessary for meticulous perceptual tasks.
Secondly, the nature of MLLM pre-training data and their resulting operational strengths may not fully align with the demands of certain fundamental perception challenges. MLLMs are typically pre-trained on vast web-scale corpora where images are often paired with textual descriptions or interleaved image-text documents Liu et al. (2023, 2024a); Agrawal et al. (2024); Bai et al. (2023a, 2025); Achiam et al. (2023). Such datasets, while fostering robust semantic alignment and impressive contextual understanding, might underrepresent stimuli requiring intensive perceptual organization purely from low-level visual features, without strong linguistic or clear object-based anchors. Consequently, MLLMs may not have adequately developed the specialized visual routines needed for tasks like consistently grouping spatially distributed elements based only on subtle shared properties (e.g., color similarity) within a field of similar distractors. Their celebrated success in many multimodal tasks often stems from leveraging their powerful integrated Large Language Models (LLMs) to interpret, reason about, and generate language from visual input Achiam et al. (2023); Anthropic (2025); Liu et al. (2023); Gong et al. (2023); Jiang et al. (2023). This reliance on higher-level, often text-mediated, understanding can be less effective when the core challenge demands direct, bottom-up visual processing and pattern extraction, rather than conceptual inference or semantic association. Such tasks require a foundational visual acuity that may not be a primary emergent outcome of training regimes largely focused on multimodal semantics and instruction following.
6 Conclusion and Future Directions
In this work, we introduced HueManity, a new benchmark dataset comprising 83,850 Ishihara-style images designed to rigorously evaluate the fine-grained visual perception of Multimodal Large Language Models (MLLMs). Our comprehensive evaluations of nine MLLMs revealed a stark performance gap: these advanced models struggled significantly (e.g.,  3-34% accuracy) on tasks readily solved by human volunteers (95-100% accuracy) and traditional computer vision models like a fine-tuned ResNet50 (94-96% accuracy). This highlights a critical limitation in current MLLM capabilities to discern patterns based on subtle visual cues in complex, noisy environments. We provide the dataset and generation code to the community to foster further research.
The observed MLLM deficiencies likely stem from a combination of factors, including vision encoders not optimized for such fine-grained detail, information bottlenecks at the vision-language interface, pre-training data that underrepresents such perceptual challenges, and an over-reliance on text-mediated or high-level semantic reasoning for tasks demanding direct, bottom-up visual processing. Future work should therefore focus on several key directions to bridge this perceptual gap: 1) Developing novel MLLM architectures with improved visual front-ends and fusion mechanisms that better preserve and process low-level visual information. 2) Augmenting pre-training and fine-tuning datasets with stimuli specifically designed to enhance robust perception in visually complex scenarios. 3) Exploring training objectives that explicitly encourage the development of fundamental visual acuity and perceptual organization skills, independent of, yet complementary to, high-level reasoning. Addressing these areas will be crucial for advancing MLLMs towards more holistic and human-like visual understanding.
7 Limitations
While HueManity provides valuable insights into a specific facet of MLLM visual perception, we acknowledge certain limitations. Firstly, the benchmark evaluates a focused perceptual task: identifying two-character alphanumeric strings within Ishihara-style dot patterns. Although this serves as a robust proxy for the ability to discern patterns from visually complex backgrounds based on subtle cues, its direct generalization to the full spectrum of diverse, real-world fine-grained perceptual challenges MLLMs may encounter requires further investigation. Secondly, our current dataset, while substantial and systematically generated with controlled color contrasts, primarily explores variations in color and basic character forms. Future iterations or complementary benchmarks could expand to assess a wider array of perceptual dimensions, such as sensitivity to texture, orientation, motion, or more intricate compositions of embedded elements. Finally, while our study includes a diverse set of nine contemporary MLLMs, the MLLM landscape is evolving rapidly. Consequently, our findings represent a snapshot based on the models and their versions tested at the time of this study, and newer or differently architected models might exhibit different performance characteristics.
8 Usage of Generative AI tools
We utilized Generative AI tools to help improve the language, phrasing, and readability of this manuscript.
Appendix A A Brief Discussion on CIEDE2000 Color Difference
$\Delta E_{2000}$ $=\sqrt{\left(\frac{\Delta L^{\prime}}{K_{L}S_{L}}\right)^{2}+\left(\frac{\Delta C^{\prime}}{K_{C}S_{C}}\right)^{2}+\left(\frac{\Delta H^{\prime}}{K_{H}S_{H}}\right)^{2}+R_{T}\left(\frac{\Delta C^{\prime}}{K_{C}S_{C}}\right)\left(\frac{\Delta H^{\prime}}{K_{H}S_{H}}\right)}$
$\text{where}\quad&\Delta L^{\prime}\text{ is the corrected lightness difference,}\\ &\Delta C^{\prime}\text{ is the corrected chroma difference,}\\ &\Delta H^{\prime}\text{ is the corrected hue difference,}\\ &K_{L},K_{C},K_{H}\text{ are parametric factors (typically 1),}\\ &S_{L},S_{C},S_{H}\text{ are weighting functions for lightness, chroma, and hue,}\\ &R_{T}\text{ is a rotation term accounting for hue-chroma interaction.}\end{aligned}$ (1)
The CIEDE2000 score ($\Delta E_{2000}$, Equation A) Luo et al. (2001) quantifies the perceived difference between two colors more accurately than prior formulae, especially for subtle variations. It calculates a single value representing the “distance” between colors in the perceptually uniform CIE $L^{*}a^{*}b^{*}$ space, considering lightness, chroma, and hue. In the HueManity benchmark, $\Delta E_{2000}$ was pivotal for systematically designing stimuli. The ability to discern characters in the Ishihara-style plates directly depends on the perceived color contrast between foreground (character) and background dots. This score provided a perceptually relevant, objective method to quantify this contrast, enabling the selection of color pairs across a controlled spectrum of difficulty, refer to Figure 3. This ensured the benchmark could rigorously test visual perception for varying degrees of color discriminability while maintaining stimuli legibility for human comparison, forming a foundational aspect of our dataset’s controlled experimental design.
Figure 3: Distribution of CIEDE2000 color difference scores for the 25 selected foreground-background color pairs utilized in the HueManity benchmark.
Appendix B Color Pairs Selection
The selection of appropriate color pairs for the foreground (characters) and background dots was a critical phase in the development of HueManity, undertaken with considerable care to ensure a balance between perceptual challenge and unambiguous human legibility. The process involved several stages:
  1. 1.
    Initial Candidate Generation: We bootstrapped the process with  15 medium-contrast color pairs generated by LLMs (Gemini, ChatGPT). This initial pool was iteratively refined by evaluating pairs against CIEDE2000 ($\Delta E_{2000}$, Eq. A) scores and visual checks. We retained promising candidates, modified some, and discarded others, while simultaneously manually crafting and vetting new pairs to meet the benchmark’s final requirements (detailed below). This refinement cycle culminated in the selection of 25 distinct pairs for the subsequent validation stages.
  2. 2.
    Quantitative Contrast Filtering (CIEDE2000): Each of these candidate pairs then underwent rigorous quantitative analysis using the CIEDE2000 ($\Delta E_{2000}$) color difference formula (Equation A). This formula is a standard measure in color science, designed to reflect perceptually meaningful differences as perceived by humans. We established a specific target range for the $\Delta E_{2000}$ score, retaining only pairs with contrast values between 25 and 75. The lower bound of 25 was set to ensure sufficient theoretical distinguishability for individuals with normal color vision, preventing pairs that would be inherently ambiguous. The upper bound of 75 aimed to exclude pairs with excessively high contrast, which might render the perceptual task trivial and deviate from the subtle challenge intended.
  3. 3.
    Balanced Contrast Distribution: A key objective during selection was to ensure the benchmark included stimuli across a spectrum of difficulty levels related to color similarity. Therefore, we deliberately curated the final set of 25 color pairs to achieve an approximately equal distribution around a $\Delta E_{2000}$ score of 50. This threshold is grounded in color science literature, often considered a point distinguishing more subtle (scores <50) from more clearly distinct (scores >50) color differences. We aimed for roughly half the selected pairs to fall below this threshold and half above, ensuring HueManity evaluates performance across varying, literature-informed degrees of color contrast difficulty.
  4. 4.
    Manual Verification and Legibility Check: Recognizing that a single numerical contrast score like $\Delta E_{2000}$ captures overall perceived difference but may not fully account for the complex interplay of hue, saturation, and luminance components, especially when rendered as dots and subjected to further transformations (gradient, color, and light shifts as described in Section 3.2), a crucial final step of manual verification was performed. As you noted, it is hard to quantify the nuanced visual impact of these combined factors with a single metric. Therefore, for every color pair that passed the quantitative filtering, sample HueManity images were generated. These renderings were meticulously inspected by the authors. The primary goal was to reject pairs where the characters, despite an acceptable overall contrast score, appeared visually too similar to the background due to the specific combination of hue, saturation, luminance, or the effect of the applied shifts. This ensured that the embedded alphanumeric characters were clearly legible and that the pattern recognition was unambiguous for human observers with normal color vision. Any pairs that resulted in ambiguous characters or were otherwise problematic during this visual check were discarded.
This multi-stage process, combining LLM-based idea generation, principled quantitative filtering based on color science, a balanced distributional strategy, and crucial human judgment to account for complex visual interactions, resulted in the final curated set of 25 color pairs. This ensures that the stimuli used in HueManity are not only theoretically sound but also practically validated for fairness, legibility, and the intended level of perceptual challenge.
Appendix C Qualitative Analysis of MLLM Failure Patterns
This section details common failure patterns observed in Multimodal Large Language Models (MLLMs) when tasked with identifying alphanumeric characters embedded in Ishihara-style dot patterns from the HueManity dataset. These observations stem from a comparative analysis of MLLM responses against human performance and ground truth data. Notably, human visual perception proved highly accurate on these tasks, with any infrequent errors typically involving confusion between graphically similar characters. In contrast, MLLMs exhibited distinct and more fundamental failure modes.
C.1 Prevalent Hallucination of Unrelated or Overly Complex Characters
A dominant failure mode across multiple MLLMs was the generation of characters, words, or even entire phrases that bore no resemblance to the two-character ground truth. This phenomenon of “hallucination” often resulted in outputs significantly more complex or contextually incongruous than the target stimuli. For instance, in the alphanumeric task, a model such as Claude 3.7 Sonnet might interpret a simple two-letter combination as a short phrase (e.g., responding with “MUST SEE” or “SOLU” for simple targets like “Rw” or “Tv”). Similarly, llava-7b could produce non-sensical strings like “HQJHSTOS”, and LLaVA-13b occasionally generated contextually unrelated phrases like “[G3T1NGST4RT3D]”. The numeric task was not immune — for a two-digit number, Claude 3.7 Sonnet was observed to list a sequence of unrelated two-digit numbers. This pattern suggests that when the fine-grained perceptual challenge overwhelms the MLLMs’ visual processing, they may default to generating text that, while perhaps linguistically plausible, is detached from the actual visual content.
C.2 Frequent Resort to Descriptive Evasion or Explicit Admission of Inability
Rather than consistently attempting to identify the embedded characters, many MLLMs frequently defaulted to one of two evasive strategies: providing a general description of the image (often correctly identifying it as a color vision test) or explicitly stating their incapacity to discern any characters. This behavior contrasted significantly with human participants, who invariably attempted the identification task. For example, models like GPT-4.1 Mini and Mistral-small3.1-24b, when presented with alphanumeric stimuli, often responded by describing the image as an Ishihara test but stated they could not clearly identify specific characters. In the numeric task, Claude 3.7 Sonnet sometimes offered similar descriptive evasions, asserting no number was visible and describing the circular dot pattern. Furthermore, some models, such as LLaVA-34b, occasionally provided categorical statements of inability, indicating they could not recognize or interpret images and requesting a description or textual input instead. This pattern suggests that MLLMs may possess internal confidence thresholds that, when triggered by low-confidence visual parsing, lead to evasive or pre-programmed “unable to process” responses rather than a forced, best-guess attempt at character recognition.
C.3 Erratic, Unpredictable, and Systematically Flawed Output Patterns
MLLM outputs were frequently characterized by their erratic and unpredictable nature. This included the generation of seemingly random strings of characters, peculiar systematic but incorrect patterns, or extreme numerical inventions far removed from the two-character target. This high variance in error types was observed both across different models for the same input and within the outputs of a single model across different images. For instance, when presented with the same alphanumeric target (e.g., “Wh”), while one model (GPT-4.1) might respond almost correctly, others exhibited diverse failures: Claude 3.7 Sonnet produced an unrelated number (“4726”); LLaVA-13b generated an exceptionally long string of sequential numbers; and Qwen VL Max incorrectly reasoned the presence of a different number (“12”). Some incorrect outputs also suggested flawed systematic processing, such as LLaVA-13b responding with a patterned string like “[L1L1L1]” for one target or generating extremely long, patterned numeric strings for others. Lengthy, seemingly gibberish character strings were also common from models like LLaVA-7b. This unpredictability underscores a lack of robust and stable visual feature extraction and interpretation, contrasting with human visual processing, which tends towards predictable errors based on similarity.

### v5

@@
-HueManity: Probing Fine-Grained Visual Perception in MLLMs
+
 Abstract
-Multimodal Large Language Models (MLLMs) excel at high-level visual reasoning, but their performance on nuanced perceptual tasks remains surprisingly limited.
-We present HueManity, a benchmark designed to assess visual perception in MLLMs.
-The dataset comprises 83,850 images featuring two-character alphanumeric strings embedded in Ishihara test style dot patterns, challenging models on precise pattern recognition.
-Our evaluation of nine state-of-the-art MLLMs on HueManity demonstrates a significant performance deficit compared to human and traditional computer vision baselines.
-The best-performing MLLM achieved a 33.6% accuracy on the numeric ‘easy’ task and a striking 3% on the alphanumeric ‘hard’ task.
-In contrast, human participants achieved near-perfect scores (100% and 95.6%) [^1] Human evaluations utilized 100-image representative subsets, sampled from the model evaluation sets for each respective task., and a fine-tuned ResNet50 model reached accuracies of 96.5% and 94.5%.
-These results highlight a critical gap in the visual capabilities of current MLLMs.
-Our analysis further explores potential architectural and training-paradigm factors contributing to this perceptual gap in MLLMs.
-We open-source HueManity dataset and code to foster further research in improving perceptual robustness of MLLMs.
-Code: https://github.com/rynaa/huemanity Dataset: https://huggingface.co/datasets/Jayant-Sravan/HueManity
-HueManity: Probing Fine-Grained Visual Perception in MLLMs
-Rynaa Groverthanks: Equal Contribution Google rynaa@google.com                                   Jayant Sravan Tamarapalli11footnotemark: 1 Google jayantsravan@google.com
-Sahiti Yerramilli11footnotemark: 1 Google sahitiy@google.com                                   Nilay Pande11footnotemark: 1 Waymo nilayp@waymo.com
-Figure 1: HueManity — A new benchmark for MLLM fine-grained visual perception.
-The pipeline creates a character mask from alphanumeric characters and renders it as an Ishihara-style pattern.
-While models achieve high accuracy on clear masks, they struggle significantly on the challenging pattern images.
+Recent Multimodal Large Language Models (MLLMs) demonstrate strong high-level visual reasoning on tasks such as visual question answering and image captioning.
+Yet existing benchmarks largely overlook their ability to capture fine-grained perceptual details.
+As MLLMs are increasingly deployed in safety and reliability critical settings, perceptual acuity becomes essential.
+We present HueManity, a scalable automated benchmark for assessing fine-grained visual perception in MLLMs.
+HueManity comprises 83,850 Ishihara-style images embedding alphanumeric strings, designed to evaluate pattern recognition, a core aspect of visual understanding.
+Our evaluation of nine state-of-the-art MLLMs uncovers a striking performance deficit: the strongest model achieved only 33.6% accuracy on a simple numeric task and 3% on a harder alphanumeric task, compared to near-ceiling performance from humans (99.38%, 93.25%) and a fine-tuned ResNet-50 (96.5%, 94.5%).
+These findings expose a critical weakness in MLLMs’ perceptual grounding, one that remains obscured by conventional benchmarks emphasizing high-level semantics.
+Code: https://github.com/rynaa/huemanity
+Dataset: https://huggingface.co/datasets/Jayant-Sravan/HueManity
+Figure 1: We present HueManity - an automated scalable benchmark for evaluation of fine-grained visual perception in MLLMs.
+The pipeline (left) embeds characters within challenging Ishihara-style image patterns, while ensuring human readability of the generated images.
+Experiments reveal (right) that humans and a fine-tuned ResNet50 baseline significantly outperform top-5 leading closed-source and open-source MLLMs, exposing a critical lack of fine-grained understanding.
 1 Introduction
-The trajectory of Multimodal Large Language Models (MLLMs) Team et al.
-(2023); Achiam et al.
-(2023); Bai et al.
-(2023a); Li et al.
-(2023b); Gong et al.
-(2023); Liu et al.
-(2024a, 2023); Anthropic (2025) in recent years has been marked by impressive advancements, demonstrating sophisticated capabilities in bridging visual and textual information.
-Their capabilities extend well beyond simple image labeling Russakovsky et al.
-(2015); Deng (2012), enabling complex tasks like generating detailed image descriptions Dong et al.
-(2024); Fu et al.
-(2024a), answering intricate visual questions requiring inference about relationships and activities Weng et al.
-(2025); Chen et al.
-(2025); Kuang et al.
-(2024), and participating in nuanced dialogue about visual content Cao et al.
-(2024).
-A key factor in their success is pre-training on vast web-scale image-text datasets, which facilitates learning powerful representations that capture high-level semantic links between visual features and language Jia et al.
-(2021); Radford et al.
-(2021); Schuhmann et al.
-(2022); Alayrac et al.
-(2022); Qi et al.
-(2020); Zhai et al.
-(2022); Pham et al.
-(2023).
-Consequently, MLLMs perform strongly on tasks benefiting from this conceptual understanding, including recognizing common objects, interpreting general scene structure, and relating them to text.
-However, the predominant evaluation paradigms for MLLMs have centered on these conceptual capabilities, largely overlooking their fine-grained perceptual acuity Bai et al.
-(2023b); Li et al.
-(2024, 2023a); Xu et al.
-(2024); Yin et al.
-(2023); Liu et al.
-(2023).
-This leaves a critical gap in understanding their performance on tasks demanding precise visual discernment – such as intricate pattern recognition, subtle feature differentiation, and robust segregation within visually cluttered background.
-Unlike tasks solvable through broad semantic association, these perceptual challenges require a more fundamental visual processing capacity, akin to the human ability to resolve intricate details.
-This paper introduces a benchmark specifically designed to probe this nuanced dimension of MLLM performance.
-Our methodology draws inspiration from the principles of Ishihara plates Clark (1924), a technique traditionally employed in human ophthalmology to assess color vision by embedding figures (like numbers or paths) within fields of multicolored, varied-size dots.
+The trajectory of Multimodal Large Language Models (MLLMs) [13, 1, 5, 27, 19, 29, 4] has been marked by impressive advancements, demonstrating sophisticated capabilities in bridging visual and textual information.
+Their skills extend beyond simple image labeling [36, 14], enabling complex tasks like generating detailed image descriptions [15, 18], answering intricate visual questions requiring inference about relationships and activities [42, 10, 24], and participating in nuanced dialogue about visual content [9].
+This success is largely attributed to pre-training on vast, web-scale image-text datasets, which has cultivated a powerful ability to map high-level semantic concepts between modalities [23, 34, 38, 3, 33, 46, 32].
+However, this focus on semantic understanding, both in training and in predominant evaluation paradigms [7, 25, 26, 43, 45, 29], has created a critical blind spot: the models’ foundational perceptual acuity remains largely unprobed.
+Human vision is not merely a semantic engine, it is fundamentally a system for perceptual organization, excelling at extracting coherent signals from visually cluttered environments.
+This ability to discern patterns from subtle, low-level cues — like color, luminance, and texture — is a prerequisite for virtually all higher-level visual reasoning.
+Existing benchmarks have predominantly centered on the conceptual capabilities of MLLMs, leaving their resilience to perceptual challenges like pattern recognition and feature differentiation in cluttered scenes largely unevaluated.
+This paper introduces HueManity, a benchmark specifically designed to probe this gap.
+Our methodology is inspired by Ishihara plates [12], a classic tool from human ophthalmology designed to isolate a specific perceptual skill: figure-ground segregation based on subtle color cues.
 It is crucial to clarify that HueManity does not aim to diagnose ‘color blindness’ in MLLMs.
-Instead, our Ishihara-style stimuli, created using controlled generation techniques, rigorously test an MLLM’s fundamental ability to identify embedded alphanumeric characters by their subtle color and luminance contrasts within visually cluttered dot patterns.
-Success on the HueManity benchmark serves as a crucial indicator of an MLLM’s potential for robust visual understanding in complex, real-world scenarios.
-Unlike often curated benchmark datasets, real-world visual environments are frequently characterized by clutter, partial occlusions, variable lighting, and unconventional information presentation.
-The ability of an MLLM to reliably parse characters in our Ishihara-style plates is intended to assess its resilience to visual clutter and its pattern recognition capabilities — foundational skills often linked to dependable performance in challenging visual settings.
-Furthermore, mastering this level of fine-grained perception is foundational for more intricate forms of visual reasoning.
-An MLLM’s struggles with basic pattern recognition under challenging conditions could imply difficulties in performing higher-order tasks that depend on the accurate interpretation of subtle details.
-Thus, HueManity serves not merely as a test of pattern recognition, but as a probe into architectural or training limitations that hinder MLLMs from achieving comprehensive, detailed visual intelligence.
-To address this identified gap and facilitate further research in this domain, this paper makes the following specific contributions:
+Instead, our Ishihara-style stimuli, created using controlled generation techniques, test an MLLM’s fundamental ability to identify embedded alphanumeric characters by their subtle color and luminance contrasts within visually cluttered dot patterns.
+The ability to parse characters from our Ishihara-style plates is a direct proxy for an MLLM’s capacity to handle real-world visual challenges characterized by clutter, partial occlusions, and variable lighting.
+An MLLM that cannot distinguish a number embedded in a field of dots may likewise struggle to perform optical character recognition (OCR) on text seen through a rainy window or accurately transcribe text from a crumpled, faded receipt.
+Therefore, HueManity serves not merely as a test of pattern recognition, but as a diagnostic tool to probe the limitations that prevent MLLMs from achieving robust, human-like visual intelligence.
+Our findings reveal that this novel evaluation exposes a profound deficit in current models, a result with significant implications for deploying MLLMs in domains requiring human-like perceptual understanding, from autonomous systems to medical imaging [44].
+To address this identified gap and facilitate further research in this domain, this paper makes the following contributions:
   1. 1.
@@
   2. 2.
-    We conducted a comprehensive evaluation of nine state-of-the-art MLLMs on HueManity’s easy and hard sets.
-Findings reveal a significant performance gap compared to strong human and fine-tuned ResNet50 baselines.
-Baseline success suggests MLLM limitations are due to their current architectures or training for fine-grained perception, rather than the task being intractable.
+    We conduct a comprehensive evaluation of nine state-of-the-art MLLMs, revealing a significant performance gap when compared to strong human and fine-tuned ResNet50 baselines.
+This suggests MLLM limitations are architectural rather than the task being intractable
   3. 3.
-    We release open-source code for generating challenging Ishihara-style perceptual stimuli (alphanumeric characters embedded in dot patterns), enabling reproducible research and community-driven extensions.
+    We release open-source code for generating challenging Ishihara-style perceptual stimuli, enabling reproducible research and community-driven extensions.
 2 Related Works
 2.1 Multimodal Models
-With the remarkable advancements of Large Language Models (LLMs), recent research has extended their capabilities to multimodal domains by integrating visual information, giving rise to Multimodal Large Language Models (MLLMs) Team et al.
-(2023); Achiam et al.
-(2023); Bai et al.
-(2023a); Li et al.
-(2023b); Gong et al.
-(2023); Liu et al.
-(2024a, 2023).
+With the remarkable advancements of Large Language Models (LLMs), recent research has extended their capabilities to multimodal domains by integrating visual information, giving rise to Multimodal Large Language Models (MLLMs) [13, 1, 5, 27, 19, 29].
 These models typically align visual features from pre-trained image encoders with LLMs via modality adaptation layers.
-Early works like BLIP-2 Li et al.
-(2023b) pioneered this architecture by first pre-training on image-text datasets and fine-tuning on task-specific benchmarks such as Visual Question Answering (VQA).
-Subsequent models like LLaVA Liu et al.
-(2023) advanced this approach by leveraging synthetic instruction-following data in VQA formats, significantly improving instruction tuning performance.
-More recent efforts have expanded into video understanding and even image generation, showcasing the versatility of MLLMs across modalities.
-However, much of their celebrated success in visual tasks often appears intertwined with, and perhaps reliant upon, their powerful language capabilities to interpret, reason about, and generate text from visual content.
-This inherent strength in linguistic processing may have led to less emphasis on developing or scrutinizing their more fundamental, non-linguistic visual perception skills.
-HueManity directly addresses this gap in scrutinizing MLLM visual perception.
-It is a benchmark specifically designed to isolate and rigorously evaluate these core abilities.
+Early works like BLIP-2 [27] pioneered this architecture by first pre-training on image-text datasets and fine-tuning on task-specific benchmarks such as Visual Question Answering (VQA).
+Subsequent models like LLaVA [29] advanced this approach by leveraging synthetic instruction-following data in VQA formats, significantly improving instruction tuning performance.
+More recent efforts have expanded into video understanding and even image generation [8, 37], showcasing the versatility of MLLMs across modalities.
+However, this celebrated success in visual tasks often appears reliant on their powerful language capabilities for reasoning and interpretation, potentially overshadowing the need to scrutinize their fundamental perception skills.
+Addressing this gap, HueManity is a benchmark designed specifically to isolate and probe these visual abilities.
 2.2 MLLM Evaluation
-Multimodal Large Language Models (MLLMs) are highly capable of global image understanding and reasoning but consistently struggle with fine-grained visual tasks such as precise recognition and localization Huang and Zhang (2024); Li et al.
-(2024).
-Several benchmarks have been proposed to address these challenges.
-TouchStone Bai et al.
-(2023b) offers 908 manually annotated visual dialog questions across five abilities and 27 sub-tasks, while LLaVA-Bench Liu et al.
-(2023) includes 24 images with 60 curated questions covering diverse content like scenes, memes, and sketches.
-To automate evaluation, LLaVA-Bench Liu et al.
-(2023), LAMM Yin et al.
-(2023), and TouchStone Bai et al.
-(2023b) rely on GPT-based models to judge relevance and accuracy, but this introduces inherent reliability issues and cost inefficiencies.
-Similarly, LVLM-eHub Xu et al.
-(2024) aggregates multiple vision benchmarks but still depends on human annotators to compare model outputs, making it both subjective and expensive.
-More structured efforts like MME and MMBench Liu et al.
-(2024b) attempt to provide objective evaluations with multiple-choice questions across a wide range of ability dimensions.
-MME pioneered multimodal Yes/No questions for perception and reasoning, with MM-Vet Yu et al.
-(2023) and MMBench Liu et al.
-(2024b) extending coverage to sub-tasks like OCR, math, and recognition, yet they all heavily rely on existing VQA datasets or GPT-generated questions.
-SEED-Bench Li et al.
-(2024, 2023a) scales up to 24,000 human-annotated multiple-choice questions covering varied input-output modalities, but its challenges remain simple, with most open-source models reaching 30–60% accuracy at the easiest level.
-Blink Fu et al.
-(2024b) attempts a more holistic evaluation across 14 perception tasks with 3,900 questions and 7,300 images, but does not assess models on combinations of tasks, and remains less challenging, as evidenced by higher model accuracies.
-HueManity uniquely evaluates Multimodal Large Language Models (MLLMs) on the foundational visual skill of discerning patterns from subtle cues in cluttered environments, employing alphanumeric characters within Ishihara-style dot patterns with controlled color contrasts as a robust proxy.
-This stimulus design, combined with procedural generation and exact-match evaluation, establishes a scalable, objective, and reliable methodology distinct from more subjective or resource-intensive techniques.
+While MLLMs excel at global image understanding, they often struggle with fine-grained tasks requiring precise recognition and localization [22, 25].
+Existing benchmarks designed to probe these limitations face three primary methodological challenges:
+Semantic Entanglement:In many perception benchmarks, like Eyes Wide Shut [40], MERLIM [41]), perception is often “entangled” with reasoning.
+Questions focusing on semantic concepts like “left/right” or “yellow animal” allow models to rely on language priors rather than pure visual input.
+Evaluation: Early benchmarks like LLaVA-Bench [29], and TouchStone [7] rely on GPT-based judges.
+However, LLM-based evaluation is unreliable for fine-grained perception since the best models are also not very adept.
+Other efforts, such as LVLM-eHub [43], depend on expensive human annotators, limiting scalability.
+Data Integrity and Scale: To improve reliability, MME [17], MMBench [30], and SEED-Bench [25] adopted objective, multiple-choice formats.
+However, these often draw from existing VQA datasets, raising concerns about data contamination.
+While ZeroBench [35] offers increased difficulty, its utility for fine-grained perception is limited by its broad focus and small scale.
+Similarly, illusion-based benchmarks [20, 39] are often constrained by small dataset sizes and a reliance on common internet images.
+HueManity departs from these approaches by explicitly disentangling perception from semantics.
+Unlike benchmarks where linguistic context could help guess an answer, HueManity provides no semantic cues to predict a hidden digit like ‘26’.
+The model must perceive the underlying dot patterns directly, making it impossible to rely on language priors.
+Through procedural generation and an exact-match evaluation framework, HueManity offers a novel, scalable, and entirely objective methodology for evaluating a core perceptual capability that remains a challenge for state-of-the-art MLLMs.
+Table 1: Comparison of MLLM evaluation benchmarks across key methodological attributes. Novelty refers to the use of non-internet, generated images.
+Benchmark Data Size Automatic Annotation Novel Images Answer Form Automated Evaluation
+LVLM-eHub - Free-form
+MMBench 3,217 Multi-choice
+SEED-Bench2 24,371 Multi-choice
+Blink 3,807 Multi-choice
+ZeroBench 100 Free-form
+HueManity 83,850 Exact Match
 3 Data Creation
-The HueManity dataset comprises 83,850 images, each pairing a two-character alphanumeric string with its ground truth label and full generation parameters.
-These strings are formed from lowercase letters (a-z), uppercase letters (A-Z), and digits (0-9).
-To enhance clarity and prevent evaluation errors, we excluded visually ambiguous characters (‘l’, ‘I’, ‘J’, ‘O’) and combinations commencing with ‘0’ (e.g., ‘01’, to avoid ‘1’ vs.
-‘01’ prediction conflicts).
-The entire dataset was produced by rendering all valid two-character combinations using 25 meticulously curated color pairs, which then formed the basis for our evaluation sets.
-The full HueManity dataset, encompassing all 83,850 generated images, serves as a comprehensive resource for evaluating MLLM perception.
-However, for the specific MLLM evaluations reported in this paper, particularly considering practical constraints such as time and API access costs, we defined two distinct tasks using carefully sampled subsets.
-Each subset comprises 1,000 images, randomly selected to be representative for these focused evaluations:
-  1. 1.
-    Number Recognition Set (Used for the ‘Easier’ Task): This subset of 1,000 images is used to evaluate models on the Number Identification task.
-It exclusively contains images displaying two-character strings composed solely of numeric digits (e.g., ‘17’, ‘83’, ‘05’).
-  2. 2.
-    Text Recognition Set (Used for the ‘Harder’ Task): This subset of 1,000 images, drawn from the broader pool of alphanumeric combinations, is employed for the Text Identification task.
-It features diverse two-character alphanumeric strings, including various combinations of letters and digits (e.g., ‘A7’, ‘b9’, ‘XG’).
-3.1 Text Mask Generator
-The initial stage in our data pipeline is the creation of a 900x900 pixel resolution binary text mask for each two-character string.
-Using Pygame, we render these strings in white against a black background.
-To ensure both clear readability and adequate character width suitable for the subsequent dot-based rendering, all masks utilize the DejaVu Sans font (size 550, bold and italic styles), as exemplified in Figure 2(a).
-(a) Binary Text Mask
-(b) Ishihara pattern
-Figure 2: Image generation process for the alphanumeric string “Y8”. (a) The binary text mask. (b) The corresponding Ishihara-style dot pattern.
-3.2 Ishihara-Style Pattern Generation
+The HueManity benchmark is built upon a dataset of 83,850 images, each featuring a two-character alphanumeric string rendered as an Ishihara-style dot pattern, along with its corresponding ground truth label and generation parameters.
+Our generation process involves creating a base text mask, selecting a color palette, and rendering the final dot pattern.
+Figure 1 illustrates the core stages of this pipeline.
+Text Mask Generator.
+The first step is to create a 900x900 pixel binary mask for each two-character string.
+We use the Pygame library to render the string in white on a black background.
+To ensure the characters are thick enough for dot-based rendering and remain clearly legible, we use the DejaVu Sans font, styled in bold and italic at a size of 550 (Figure 1).
+Color Pairs Selection To generate the patterns, we meticulously selected 25 distinct foreground-background color pairs.
+This selection process involved a multi-stage procedure combining quantitative analysis using the CIEDE2000 color-difference formula [31] with extensive manual verification.
+This ensures that the color pairs are balanced for perceptual difficulty while remaining legible to humans (see Appendix D for details).
+Ishihara-Style Pattern Generation.
 Our pattern generator, adapted from an open-source Pygame project [^1] https://github.com/hakrackete/Ishihara-color-plate-generator, iteratively populates the image with non-overlapping circles.
-Over 30,000 iterations, the generator performs the following steps for each potential circle:
-  1. 1.
-    Placement: Randomly sample (x,y) coordinates and compute the maximum non-colliding radius (ranging from 4 to 15 pixels).
-  2. 2.
-    Initial Color: Assign initial color (foreground/background) based on whether the center falls within the character mask (Fig. 2(a)).
-  3. 3.
-    Color Transformation: Apply three randomized transformations to the initial color: gradient shift towards the other pair color, RGB color shift (range $[-30,+30]$), and RGB lightness scaling (factor 0.66-1.5).
-  4. 4.
-    Rendering: Render the transformed circle at the computed position.
-This iterative process ultimately yields the final, dense Ishihara-style pattern (as shown in Figure 2(b)).
-3.3 Color Pairs Selection
-The 25 distinct foreground-background color pairs used in the HueManity stimuli were meticulously selected through a multi-stage process to ensure a balance between perceptual challenge and unambiguous human legibility.
-This procedure involved quantitative CIEDE2000 Luo et al.
-(2001) analysis and extensive manual verification, with the full methodology detailed in Appendix B.
-4 Experiments
-We evaluated a diverse set of nine state-of-the-art Multimodal Large Language Models (MLLMs), encompassing prominent commercial APIs and publicly available open-source models.
+Over 30,000 iterations, the generator randomly places circles, computes their maximum non-colliding radius (4-15 pixels), and assigns it a color.
+The color assignment depends on whether the circle’s center falls within the character region of the text mask.
+This initial color then undergoes three randomized transformations: a gradient shift towards the other color, an RGB color shift (range [-30, +30]), and an RGB lightness scaling (factor 0.66 - 1.5).
+These transformed circles are then rendered, resulting in the final dense Ishihara-style pattern (Figure 2).
+Figure 2: Qualitative examples showing predictions of 4 representative MLLMs vs baselines on numeric and alphanumeric tasks.
+4 Experimental Setup
+Benchmark Tasks and Evaluation Sets.
+From the generated stimuli, we designed a benchmark with two recognition tasks of varying difficulty and two visual conditions to isolate specific model capabilities.
+There are two distinct recognition tasks of varying difficulty: a numeric task with a label space of 90 possible outputs and a more challenging alphanumeric task with 3,364 possible outputs (Refer to Appendix B.1).
+The larger search space makes the alphanumeric task inherently more difficult.
+  • $\circ$
+    Number Recognition Set (Easier Task): This subset contains images with two-digit numeric strings (e.g., 17, 83, 65).
+We exclude leading-zero numbers (‘00’-‘09’) to prevent ambiguity between single and double-digit interpretations.
+  • $\circ$
+    Alphanumeric Recognition Set (Harder Task): This subset includes two-character strings (e.g., A7, 9b, XG) formed from lowercase letters (a-z), uppercase letters (A-Z), and digits (0-9).
+We excluded visually ambiguous characters (‘l’, ‘I’, ‘J’, ‘O’) from the character space.
+Further, to disentangle the challenge of perceptual grouping from fundamental character recognition, we evaluate models under two visual conditions for each image.
+Due to computational and API cost constraints, our MLLM evaluations were conducted on subsets of 1,000 images, randomly sampled for both the numeric and alphanumeric tasks.
+  • $\circ$
+    Ishihara Pattern: Primary task using the Ishihara-style dot pattern images.
+  • $\circ$
+    Text Masks: Control task using the binary text mask used in the creation of Ishihara Pattern images.
+This condition establishes a baseline for each model’s fundamental OCR capability, helping to disentangle it from performance on the more perceptually complex patterns.
+Multimodal Large Language Models (MLLMs).
+We evaluated a diverse set of nine Multimodal Large Language Models (MLLMs), including both commercial API-based and publicly available open-source models.
 Model inference was orchestrated using Promptfoo [^2] https://www.promptfoo.dev, a platform facilitating reproducible benchmarking through flexible prompt definition and API integration.
 Open-source models were hosted locally via Ollama [^3] https://ollama.com and inferred on a single NVIDIA A100 GPU.
-Evaluation focused on two primary tasks: numerical recognition (digits only) and alphanumeric recognition (digits and letters).
-For each of these tasks, a subset of 1,000 points from HueManity: two-character strings and their corresponding generated images were used.
-Models were evaluated on two types of visual stimuli for each datapoint:
-  • HueManity Plates: The 1,000 randomly sampled Ishihara-style dot pattern images from the HueManity benchmark.
-  • Text Masks: The corresponding 1,000 binary text mask images (i.e., white text on a black background, as exemplified in Figure 2(a)).
-This additional evaluation on text masks was performed to establish a baseline for each model’s ability to recognize the characters in a clear, unobstructed format, helping to differentiate fundamental OCR capabilities from performance on the perceptually challenging dot patterns.
-All images (both HueManity plates and text masks) were provided to the models encoded in Base64 format. The following task-specific prompts were used for both stimulus types:
-  • Number Recognition Prompt: "What is the number in this image? Strictly stick to the format: Answer: [number in the image]"
-  • Text Recognition Prompt: "What is the exact text in this image?
+Images were Base64 encoded at 900x900 original resolution and submitted with the following task-specific prompts, which were kept consistent across all models and conditions:
+Number Recognition Prompt: “What is the number in this image?
+Strictly stick to the format: Answer: [number in the image]” Text Recognition Prompt: “What is the exact text in this image?
 It has only alpha-numeric characters excluding small l, capital O, capital I, and capital J to avoid ambiguity.
-Strictly stick to the format: Answer: [exact text in the image]"
-4.1 Human Performance Evaluations
-To establish a human performance baseline, evaluations involved three adult volunteers, all with self-reported normal color vision.
-This group size was considered adequate because identifying characters in HueManity plates is a rudimentary perceptual task for humans with normal vision, one that requires no specialized skills and typically exhibits low inter-volunteer performance variance.
-For each of the two tasks (numerical and alphanumeric recognition), these volunteers were tested on a representative subset of 100 images, randomly sampled from the corresponding 1,000-image evaluation sets used for MLLMs.
-Volunteers viewed the images at their original 900x900 pixel resolution within a Google Sheets document.
-Using the identical prompts provided to the MLLMs, they were asked to identify the embedded characters, entering their responses directly into the document.
-This methodology ensured a direct comparison of human and MLLM performance under equivalent conditions.
-4.2 Traditional Computer Vision Baseline (ResNet50)
-As a representative traditional computer vision baseline, we trained and evaluated a ResNet50 model. We utilized a ResNet50 pre-trained on ImageNet, obtained from the PyTorch vision library.
+Strictly stick to the format: Answer: [exact text in the image]”
+Human Performance Baseline.
+To establish a human performance baseline, we recruited 16 adult volunteers from diverse age groups with self-reported normal color vision.
+Each participant was evaluated on 100 Ishihara pattern images for each of the tasks (numerical and alphanumeric recognition).
+These images were randomly sampled from the 1,000-image sets used in the MLLM evaluations.
+Participants viewed the 900x900 pixel images in a Google Sheets document and provided their responses to the same prompts given to the models, ensuring a direct and methodologically consistent comparison.
+Traditional Computer Vision Baseline (ResNet50).
+As a representative traditional computer vision baseline, we trained and evaluated a ResNet50 model.
+We utilized a ResNet50 pre-trained on ImageNet, obtained from the PyTorch vision library.
 The standard classification layer of the ResNet50 was replaced with two independent classification heads.
@@
 For the purpose of fine-tuning this model, we utilized 2,000 images randomly sampled from the broader HueManity dataset, ensuring these were distinct from the final evaluation subsets.
-Training was conducted for 30 epochs using the Adam optimizer with a learning rate of $31e-$ 3.
-The loss function for training was the sum of the cross-entropy losses calculated independently for each of the two classification heads.
+Training was conducted for 30 epochs using the Adam optimizer with a learning rate of 1e31e-3.
+The loss function was the sum of the cross-entropy losses calculated independently for each of the two classification heads.
 The trained model was evaluated on the same 1,000-image subsets used for the MLLM evaluations.
 5 Results and Analysis
-5.1 Human Performance: A Near-Perfect Baseline
-Human volunteers established a crucial performance baseline, demonstrating exceptionally high accuracy and efficiency on the HueManity benchmark (detailed in Table 3).
-On the numerical recognition task, volunteers achieved a perfect 100% average accuracy, indicating that identifying two-digit strings within the dot patterns was straightforward.
-The more complex alphanumeric task also yielded a very strong average accuracy of 95.6%.
-The minor errors in the alphanumeric task were insightful: annotator feedback consistently attributed confusion not to an inability to perceive the characters, but to difficulty differentiating visually similar character forms (e.g., ‘s’, ‘c’, ‘w’ upper and lower case variants).
-This distinction is important, as it underscores that the primary perceptual challenge of extracting patterns from the dots was readily overcome by the human visual system.
-Furthermore, volunteers processed these images with remarkable efficiency, typically requiring less than one second per image to identify the characters.
-Near-perfect and rapid human scores offer crucial insights into MLLM performance on this task for several key reasons:
-  • Confirms Task Solvability and Stimulus Clarity: High human accuracy confirms the stimuli provide unambiguous visual information for character identification, showing the tasks are well-defined and solvable by proficient visual systems.
-  • Establishes a Performance Ceiling: Human performance sets a clear accuracy ceiling, highlighting the target for machine perception systems on this (for humans) perceptually simple task.
-  • Highlights the Nature of MLLM Deficiencies: Significant MLLM deviations from human scores indicate machine-specific limitations, not task impossibility or unclear stimuli.
-This contrast helps determine if MLLM struggles lie in initial perceptual grouping or later recognition.
-  • Contextualizes Machine Errors: Human ease and minor confusions (e.g., similar glyphs) provide valuable context for analyzing MLLM error types and severity, pinpointing where machine perception significantly diverges from human visual understanding.
-In essence, the human baseline shows HueManity tests foundational visual skills that humans execute efficiently, despite MLLM challenges.
-This context is vital for appreciating the scale and nature of MLLM performance gaps.
-5.2 ResNet50 Baseline: Demonstrating Task Learnability
-To further contextualize the performance of Multimodal Large Language Models (MLLMs), we established a strong baseline using a fine-tuned ResNet50 He et al.
-(2015), a well-established traditional computer vision architecture.
-As detailed in Table 1 and Table 2, the ResNet50 model achieved high accuracy, scoring 96.5% on the numerical recognition task and 94.5% on the alphanumeric task.
-This robust performance from a standard convolutional architecture, fine-tuned on a relatively small set of 2,000 images from the HueManity dataset (as described in Section 4.2), is highly informative.
-It clearly demonstrates that the task of identifying characters within the HueManity dot patterns is fundamentally learnable by established computer vision techniques.
-Furthermore, achieving near-human accuracy with this setup suggests that the perceptual cues embedded in the stimuli are sufficiently rich and consistent for a focused model to learn effective recognition.
-This implies that the task, while designed to challenge nuanced perception, is not inherently intractable for AI and can be considered relatively easy for appropriately adapted vision models.
-The ResNet50 baseline’s success, therefore, pinpoints the challenges as MLLM-specific.
-This suggests that MLLM difficulties likely arise not from an inherently unsolvable visual task, but from how these larger, more general models currently process or prioritize fine-grained perceptual information.
-Table 1: Accuracy on the number recognition task for human evaluators, ResNet50, and various MLLMs on both text masks and patterned HueManity images.
-Mask Patterned Image
-Humans (average) 100% 100%
-ResNet50 - 96.5%
+Table 2: Accuracy on the number and alphanumeric recognition tasks for human evaluators, ResNet50, and various MLLMs on both text masks and patterned HueManity images.
+$\pm$ denotes the Wilson Confidence intervals at 95% confidence.
+Number Task Alphanumeric Task
+Mask Pattern Mask Pattern
+Random choice 1.11% 1.11% 0.029% 0.029%
+Humans (average) - 99.37 $\pm$ 2.37% - 93.25 $\pm$ 5.08%
+ResNet50 - 96.5 $\pm$ 1.15% - 94.5 $\pm$ 1.42%
 API-based models
-GPT-4.1 mini 100% 19.0%
-GPT-4.1 100% 33.6%
-Claude 3.7 Sonnet 100% 0.4%
+GPT-4.1 100 $\pm$ 0.19% 33.6 $\pm$ 2.92% 80 $\pm$ 2.47% 3.0 $\pm$ 1.07%
+GPT-4.1 mini 100 $\pm$ 0.19% 19.0 $\pm$ 2.42% 72.4 $\pm$ 2.76% 0.6 $\pm$ 0.51%
+Claude 3.7 Sonnet 100 $\pm$ 0.19% 0.4 $\pm$ 0.43% 82.2 $\pm$ 2.36% 0 $\pm$ 0.19%
 Open-source models
-LLaVA-v1.6-7B 87.7% 3.3%
-LLaVA-v1.6-13B 87.2% 8.1%
-LLaVA-v1.6-34B 96.6% 7.8%
-Mistral-small3.1-24b 100% 0.1%
-Qwen VL Max 100% 0.2%
-Pixtral 100% 1%
-Table 2: Accuracy on the alphanumeric recognition task, comparing human performance with ResNet50 and MLLMs across text masks and patterned HueManity images.
-Mask Patterned Image
-Humans (average) 100% 95.6%
-ResNet50 - 94.5%
-API-based models
-GPT-4.1 mini 72.4% 0.6%
-GPT-4.1 80% 3.0%
-Claude 3.7 Sonnet 82.2% 0%
-Open-source models
-LLaVA-v1.6-7B 15% 0%
-LLaVA-v1.6-13B 31.8% 0.1%
-LLaVA-v1.6-34B 27.1% 0%
-Mistral-small3.1-24b 58.7% 0%
-Qwen VL Max 83.5% 0%
-Pixtral 65.8% 1.8%
-Table 3: Human evaluation accuracies on 100-image subsets for both number and alphanumeric pattern recognition tasks.
-Number Pattern Alphanumeric Pattern
-Evaluator 1 100% 97%
-Evaluator 2 100% 95%
-Evaluator 3 100% 95%
-Average 100% 95.6%
-5.3 The Perceptual Gap: MLLM Performance vs. Baselines on HueManity
-The performance of the nine evaluated Multimodal Large Language Models (MLLMs) on the HueManity benchmark stands in stark contrast to the near-perfect accuracies achieved by both human volunteers and the fine-tuned ResNet50 baseline He et al.
-(2015) (detailed in Table 1 and Table 2).
-Across both the numerical and alphanumeric tasks, all evaluated MLLMs exhibited a profound and consistent inability to match these baselines.
-For instance, even the best-performing MLLM achieved only 33.6% on the easier numeric task and a mere 3% on the harder alphanumeric task—figures that dramatically underscore their struggle.
-This significant and consistent underperformance, on tasks demonstrably solvable by human vision and traditional AI methodologies, signals a critical gap in the fine-grained visual perception capabilities of current MLLMs.
-Such results necessitate a deeper exploration into the potential underlying reasons for these models’ difficulties.
-Several characteristics inherent to the current design and training paradigms of many Multimodal Large Language Models (MLLMs) may contribute to their observed difficulties on tasks demanding nuanced visual perception.
-Firstly, the visual processing pipeline itself could be a factor.
-Vision encoders in MLLMs, though often powerful and pre-trained on diverse datasets (e.g., ViT variants Liu et al.
-(2023, 2024a); Agrawal et al.
-(2024); Bai et al.
-(2023a, 2025), are typically optimized for capturing semantic information and global scene context.
-This optimization focus might lead to the unintentional down-weighting or loss of extremely fine-grained local details—like subtle color shifts defining patterns within a dense, similarly featured background—that are crucial when strong semantic cues are absent.
-Furthermore, the interface used to connect these vision encoders to the language model—often a projection layer or a small MLP Liu et al.
-(2023, 2024a); Agrawal et al.
-(2024); Bai et al.
-(2023a, 2025)—necessarily transforms and condenses visual information.
-This critical step, while enabling multimodal fusion, might inadvertently act as an information bottleneck, potentially abstracting or losing the precise, high-resolution feature distinctions necessary for meticulous perceptual tasks.
-Secondly, the nature of MLLM pre-training data and their resulting operational strengths may not fully align with the demands of certain fundamental perception challenges.
-MLLMs are typically pre-trained on vast web-scale corpora where images are often paired with textual descriptions or interleaved image-text documents Liu et al.
-(2023, 2024a); Agrawal et al.
-(2024); Bai et al.
-(2023a, 2025); Achiam et al.
-(2023).
-Such datasets, while fostering robust semantic alignment and impressive contextual understanding, might underrepresent stimuli requiring intensive perceptual organization purely from low-level visual features, without strong linguistic or clear object-based anchors.
-Consequently, MLLMs may not have adequately developed the specialized visual routines needed for tasks like consistently grouping spatially distributed elements based only on subtle shared properties (e.g., color similarity) within a field of similar distractors.
-Their celebrated success in many multimodal tasks often stems from leveraging their powerful integrated Large Language Models (LLMs) to interpret, reason about, and generate language from visual input Achiam et al.
-(2023); Anthropic (2025); Liu et al.
-(2023); Gong et al.
-(2023); Jiang et al.
-(2023).
-This reliance on higher-level, often text-mediated, understanding can be less effective when the core challenge demands direct, bottom-up visual processing and pattern extraction, rather than conceptual inference or semantic association.
-Such tasks require a foundational visual acuity that may not be a primary emergent outcome of training regimes largely focused on multimodal semantics and instruction following.
+LLaVA-v1.6-34B 96.6 $\pm$ 1.13% 7.8 $\pm$ 1.66% 27.1 $\pm$ 2.75% 0 $\pm$ 0.19%
+LLaVA-v1.6-13B 87.2 $\pm$ 2.07% 8.1 $\pm$ 1.69% 31.8 $\pm$ 2.88% 0.1 $\pm$ 0.27%
+LLaVA-v1.6-7B 87.7 $\pm$ 2.03% 3.3 $\pm$ 1.11% 15 $\pm$ 2.21% 0 $\pm$ 0.19%
+Mistral-small3.1-24b 100 $\pm$ 0.19% 0.1 $\pm$ 0.27% 58.7 $\pm$ 3.04% 0 $\pm$ 0.19%
+Qwen VL Max 100 $\pm$ 0.19% 0.2 $\pm$ 0.33% 83.5 $\pm$ 2.29% 0 $\pm$ 0.19%
+Pixtral 100 $\pm$ 0.19% 1 $\pm$ 0.64% 65.8 $\pm$ 2.93% 1.8 $\pm$ 0.84%
+5.1 Human and ResNet50 Baseline Performance
+To contextualize the performance of MLLMs, we established two critical baselines: human evaluators and a fine-tuned ResNet50 model.
+These baselines serve to confirm the task’s solvability and provide a clear performance ceiling.
+Human evaluators demonstrated near-perfect accuracy, achieving 99.37% on the numerical task and 93.25% on the alphanumeric task.
+Annotator feedback indicated that the few errors on the alphanumeric task were due to confusion between visually similar characters (e.g., ‘s’ vs.
+‘S’, ‘c’ vs.
+‘C’, ‘w’ vs.
+‘W’), not an inability to perceive the patterns.
+Annotators also report that the character recognition time was in the order of a few seconds.
+This near-flawless performance confirms that the stimuli are clear and the task is fundamentally solvable for a proficient visual system.
+Similarly, a fine-tuned ResNet50 [21] model achieved high accuracy, scoring 96.5% on the numerical task and 94.5% on the alphanumeric task.
+This strong performance from a standard convolutional architecture, trained on only 2,000 examples, proves that the perceptual cues within the images are sufficient for established computer vision techniques to learn the task.
+Together, these baselines establish that the HueManity benchmark is a well-posed and learnable perception challenge, suggesting that any failures by MLLMs are likely due to architectural or training-related limitations rather than the inherent difficulty of the task itself.
+5.2 MLLM Performance
+The performance of the nine evaluated MLLMs on HueManity sharply contrasts with the near-perfect accuracies of both humans and the ResNet50 baseline (Table 2).
+All MLLMs consistently struggled, with the best achieving only 33.6% on the numeric task and a mere 3% on the alphanumeric task.
+As expected, most models demonstrated strong Optical Character Recognition (OCR) capabilities on the simple binary mask images.
+For the numeric mask task, nearly all models achieved perfect or near-perfect accuracy.
+Performance on the more complex alphanumeric mask was more varied, while API-based models like GPT-4.1 and Claude 3.7 Sonnet maintained high accuracy (>>80%), the open-source LLaVA family notably struggled (15-32% accuracy), indicating a weaker baseline OCR capability for complex character sets.
+The models’ high OCR accuracy on masks makes their widespread failure on the patterned images even more striking.
+This drop in performance pinpoints the challenge as a perceptual failure, not a simple recognition one.
+On the “easy” numeric pattern task, a clear performance hierarchy emerged.
+GPT-4.1 was the only model with meaningful success at 33.6% accuracy.
+A second tier, including GPT-4.1 mini (19.0%) and the LLaVA family (3.3-8.1%), showed minimal capability.
+The remaining models, such as Claude 3.7 Sonnet and Qwen VL Max, failed almost completely, scoring near zero.
+On the “hard” alphanumeric pattern task, MLLM performance collapsed almost entirely.
+GPT-4.1 was again the top performer but with a mere 3.0% accuracy.
+It was followed by Pixtral at 1.8%.
+All other evaluated models scored less than 1%, demonstrating a near-total inability to perform perceptual grouping on this more complex task.
+Several characteristics inherent to the current design and training paradigms of many MLLMs may contribute to their observed difficulties on tasks demanding nuanced visual perception.
+Semantic Optimization in Pre-Training.
+MLLM vision encoders are typically optimized to capture high-level semantic information for tasks like scene understanding [29, 28, 2, 5, 6].
+This focus on global context may cause the loss of fine-grained local details, such as the subtle color and texture cues that define the characters in HueManity.
+Furthermore, pre-training on web-scale datasets of images paired with descriptive text may not provide sufficient exposure to stimuli requiring intensive perceptual organization without strong semantic anchors.
+Architectural Bottlenecks.
+The architectures themselves may contribute to this failure.
+Many MLLMs use Vision Transformers (ViTs) [16] that divide images into patches, a process that can disrupt fine-grained details within those patches.
+Additionally, the projection layers that connect the vision encoder to the language model can act as an information bottleneck, abstracting away precise, high-resolution feature distinctions that are essential for solving our task.
+5.3 Can MLLMs Learn the Task? Probing with In-Context Learning and Fine-Tuning
+In-Context Learning We tested whether providing in-context examples could improve performance.
+While this approach retained 100% accuracy for the simple “mask” images, it proved ineffective and often detrimental for the patterned Ishihara plates (Table 3).
+For most models, performance degraded as more examples were added.
+These results are consistent with [39]’s findings and we hypothesize that the visual complexity of the examples introduces noise that confuses the models, reinforcing the conclusion that the deficit is perceptual rather than a lack of contextual understanding.
+Table 3: Few-shot performance on Ishihara Pattern for number task with 95% Wilson confidence intervals.
+Model 0-shot (Baseline) 1-shot 3-shot 5-shot
+GPT 4.1 33.6 $\pm$ 2.9% 34.0 $\pm$ 2.9% 35.0 $\pm$ 3.0% 31.0 $\pm$ 2.9%
+GPT-4.1 mini 19.0 $\pm$ 2.5% 14.0 $\pm$ 2.2% 9.0 $\pm$ 1.8% 6.0 $\pm$ 1.5%
+Pixtral Large 1.0 $\pm$ 0.6% 1.0 $\pm$ 0.6% 0.0 $\pm$ 0.2% 0.0 $\pm$ 0.2%
+Claude 3.7 Sonnet 0.4 $\pm$ 0.4% 0.0 $\pm$ 0.2% 1.0 $\pm$ 0.6% 1.0 $\pm$ 0.6%
+Qwen 2.5 VL Max 0.2 $\pm$ 0.3% 0.0 $\pm$ 0.2% 0.0 $\pm$ 0.2% 0.0 $\pm$ 0.2%
+Table 4: Performance of fine-tuned Gemma3-4B and qualitative examples of predictions.
+Training Data Task Accuracy Predictions for Ground Truth Values
+ML sZ cT 1Q
+N/A (Zero-Shot) Mask 80% ML SZ CT 1Q
+Pattern 0% 76492385 739482561 74928536 749238561
+Fine-Tuned on Masks Mask 94% ML sZ cT 1Q
+Pattern 0% R1 M5 4model D4
+Fine-Tuned on Ishihara Mask 0% 7s 24 n6 36
+Pattern 0% tH vF 6t 58
+MLLM Fine-tuning To determine if direct training could overcome the observed perceptual challenges, we LoRA fine-tuned the Gemma-3-4B model on 500 examples from the HueManity dataset using the HuggingFace SFTTrainer [^4] https://ai.google.dev/gemma/docs/core/huggingface_text_full_finetune with Adam optimizer, 2e-4 learning rate, and 3 epochs.
+The results of this limited experiment were striking and revealed a critical limitation.
+While fine-tuning on the simple text masks improved performance on that control task as expected, it provided no benefit for recognizing the patterned Ishihara images.
+More critically, fine-tuning on the patterned images not only failed to improve performance on the core task but also destroyed the model’s ability to solve the simple mask task, causing its accuracy to plummet from 80% to 0%.
+Instead of learning to perceive the characters, the model simply learned to output plausible-looking two-character strings (e.g., “4G”, “6t”) regardless of the visual input.
+This suggests the failure is rooted in deeper architectural limitations, as the model appears incapable of learning the necessary perceptual skill.
+5.4 Ablations and Additional Analyses
+Image Resolution Ablation.
+To understand the impact of image resolution, we conducted an ablation study by evaluating model performance across a range of resolutions from 300px to 1300px (Table 5).
+The results were twofold.
+First, models that performed poorly at the native 900px resolution, such as Qwen VL Max, failed across all tested resolutions, which reinforces the claim that they have a fundamental perceptual deficit.
+Second, we uncovered a “squinting effect” in the best-performing model, GPT-4.1, whose accuracy peaked at the lowest resolution, increasing from 34% to 49% at 300px.
+We hypothesize that downsampling acts as a low-pass filter, smoothing the high-frequency “noise” from the dot patterns and making the underlying character shapes more salient to the model.
+Table 5: Image resolution ablation with 95% Wilson confidence intervals (N=100).
+Model 300px 500px 700px 900px 1100px 1300px
+GPT 4.1 $49.0\pm 9.8\%$ $29.0\pm 8.9\%$ $31.0\pm 9.1\%$ $34.0\pm 9.3\%$ $30.0\pm 9.0\%$ $29.2\pm 8.9\%$
+GPT-4.1 mini $39.0\pm 9.6\%$ $18.2\pm 7.6\%$ $18.0\pm 7.6\%$ $18.0\pm 7.6\%$ $10.0\pm 5.9\%$ $6.0\pm 4.6\%$
+Qwen VL Max $1.0\pm 2.0\%$ $1.0\pm 2.0\%$ $1.0\pm 2.0\%$ $1.0\pm 2.0\%$ $1.0\pm 2.0\%$ $1.0\pm 2.0\%$
+Pixtral Large $1.0\pm 2.0\%$ $0.0\pm 1.9\%$ $1.0\pm 2.0\%$ $0\pm 1.9\% 0.0$$\pm 1.9\% 0.0$$0\pm 1.9\%$
+MLLM Failure Patterns.
+A qualitative analysis of MLLM failures on the HueManity benchmark reveals that, unlike humans whose errors are predictable, MLLMs exhibit three distinct and more fundamental failure patterns when their perceptual abilities are overwhelmed (see Appendix E for more details).
+  • $\circ$
+    Hallucination: Models often generate completely unrelated and overly complex text, such as turning a two-character string into a full phrase (e.g., ‘MUST SEE’).
+This happens when the perceptual challenge overwhelms their visual processing.
+  • $\circ$
+    Evasion: They frequently engage in evasive behavior, either by describing the image as an ‘Ishihara test’ without attempting to identify the characters or by explicitly stating they are unable to perform the task.
+This suggests an internal confidence threshold triggers a pre-programmed failure response.
+  • $\circ$
+    Erratic Outputs: Their outputs are often erratic and unpredictable, ranging from random strings and nonsensical numbers to flawed, repetitive patterns (e.g., ‘[L1L1L1]’).
+This unpredictability points to a lack of robust and stable visual feature extraction.
+Quantitative Correlation with Real-World Benchmarks.
+To validate that performance on HueManity translates to real-world capabilities, we correlated model rankings on our alphanumeric pattern task with the human-preference Elo ratings from Vision Arena [11], a benchmark for diverse, real-world multimodal tasks.
+Our analysis included the seven models common to both leaderboards: GPT-4.1, GPT-4.1 mini, Claude 3.7 Sonnet, Mistral Medium, Qwen VL Max, Pixtral Large, and LLaVA-v1.6-34B.
+The analysis revealed a strong and statistically significant positive correlation of $\bm{\rho=0.8214}$ ($\bm{p=0.0227}$).
+This result provides compelling empirical evidence that HueManity is not an isolated synthetic challenge but effectively measures a foundational perceptual skill that is highly predictive of an MLLM’s general performance on complex, real-world tasks.
+This validates our benchmark as a potent diagnostic tool for identifying critical and generalizable gaps in the core perceptual abilities of current models.
 6 Conclusion and Future Directions
-In this work, we introduced HueManity, a new benchmark dataset comprising 83,850 Ishihara-style images designed to rigorously evaluate the fine-grained visual perception of Multimodal Large Language Models (MLLMs).
-Our comprehensive evaluations of nine MLLMs revealed a stark performance gap: these advanced models struggled significantly (e.g.,  3-34% accuracy) on tasks readily solved by human volunteers (95-100% accuracy) and traditional computer vision models like a fine-tuned ResNet50 (94-96% accuracy).
-This highlights a critical limitation in current MLLM capabilities to discern patterns based on subtle visual cues in complex, noisy environments.
-We provide the dataset and generation code to the community to foster further research.
-The observed MLLM deficiencies likely stem from a combination of factors, including vision encoders not optimized for such fine-grained detail, information bottlenecks at the vision-language interface, pre-training data that underrepresents such perceptual challenges, and an over-reliance on text-mediated or high-level semantic reasoning for tasks demanding direct, bottom-up visual processing.
-Future work should therefore focus on several key directions to bridge this perceptual gap: 1) Developing novel MLLM architectures with improved visual front-ends and fusion mechanisms that better preserve and process low-level visual information.
-2) Augmenting pre-training and fine-tuning datasets with stimuli specifically designed to enhance robust perception in visually complex scenarios.
-3) Exploring training objectives that explicitly encourage the development of fundamental visual acuity and perceptual organization skills, independent of, yet complementary to, high-level reasoning.
-Addressing these areas will be crucial for advancing MLLMs towards more holistic and human-like visual understanding.
-7 Limitations
-While HueManity provides valuable insights into a specific facet of MLLM visual perception, we acknowledge certain limitations.
-Firstly, the benchmark evaluates a focused perceptual task: identifying two-character alphanumeric strings within Ishihara-style dot patterns.
-Although this serves as a robust proxy for the ability to discern patterns from visually complex backgrounds based on subtle cues, its direct generalization to the full spectrum of diverse, real-world fine-grained perceptual challenges MLLMs may encounter requires further investigation.
-Secondly, our current dataset, while substantial and systematically generated with controlled color contrasts, primarily explores variations in color and basic character forms.
-Future iterations or complementary benchmarks could expand to assess a wider array of perceptual dimensions, such as sensitivity to texture, orientation, motion, or more intricate compositions of embedded elements.
-Finally, while our study includes a diverse set of nine contemporary MLLMs, the MLLM landscape is evolving rapidly.
-Consequently, our findings represent a snapshot based on the models and their versions tested at the time of this study, and newer or differently architected models might exhibit different performance characteristics.
-8 Usage of Generative AI tools
-We utilized Generative AI tools to help improve the language, phrasing, and readability of this manuscript.
-Appendix A A Brief Discussion on CIEDE2000 Color Difference
+In this work, we introduce HueManity, a large-scale benchmark with 83,850 Ishihara-style images designed to probe the fine-grained visual perception of Multimodal Large Language Models (MLLMs).
+Our comprehensive evaluation of nine state-of-the-art models reveals a critical weakness: the best-performing MLLM achieved only (33.6%, 3%) accuracies on the numeric and alphanumeric Ishihara pattern recognition tasks, starkly contrasting with the near-perfect scores of human evaluators (99.37%, 93.25%) and a fine-tuned ResNet50 baseline (96.5%, 94.5%).
+We demonstrate that this perceptual deficit is not easily remedied, as neither in-context learning nor direct fine-tuning yielded improvements, suggesting the failure may be rooted in fundamental architectural limitations rather than the task’s inherent difficulty.
+The strong correlation ($\rho=0.82$) between HueManity performance and rankings on the real-world Vision Arena benchmark validates that our findings are not isolated to a synthetic task but are indicative of a broader, more generalizable weakness in MLLM perceptual abilities.
+The deficiencies observed likely stem from vision encoders optimized for high-level semantics at the expense of fine-grained detail, information bottlenecks at the vision-language interface, and pre-training datasets that lack sufficient perceptual challenges.
+To bridge this gap, future work should focus on developing novel MLLM architectures that preserve low-level visual information, augmenting training datasets with perceptually challenging stimuli, and designing training objectives that explicitly foster robust visual acuity independent of high-level semantic reasoning.
+We release our dataset and generation code to facilitate further research toward building more perceptually grounded MLLMs.
+Appendix 0.A Model Inventory
+Model Provider Version Temp / Top-p Max Tokens
+GPT-4.1 mini OpenAI gpt-4.1-mini-2025-04-14 1.0 / 1.0 1024
+GPT-4.1 OpenAI gpt-4.1-2025-04-14 1.0 / 1.0 1024
+Claude 3.7 Sonnet Anthropic claude-3-7-sonnet-20250219 1.0 / 1.0 1024
+LLaVA-v1.6-7B Ollama llava:7b 0.8 / 0.9 -1 (inf)
+LLaVA-v1.6-13B Ollama llava:13b 0.8 / 0.9 -1 (inf)
+LLaVA-v1.6-34B Ollama llava:34b 0.8 / 0.9 -1 (inf)
+Mistral-small3.1-24b Mistral AI mistral-small-3.1-24b-instruct 0.7 / 1.0 1024
+Qwen VL Max Qwen (OpenRouter) qwen/qwen-vl-max 1.0 / 1.0 1024
+Pixtral Mistral AI pixtral-large-2411 0.7 / 1.0 1024
+Table 6: Standardized evaluation settings per model.
+All models were evaluated between May 6-7, 2025.
+Ollama models were hosted on a single GPU machine with an NVIDIA A100.
+These generation values are the model/provider defaults and we stick to these values to avoid introducing noise to our evaluations.
+Appendix 0.B Label Spaces & Chance Baselines
+Character universe.
+We define a character set $\mathcal{C}$ used to compose two-character labels. Starting from decimal digits and ASCII letters with case,
+$\{0,\ldots,9,\,A,\ldots,Z,\,a,\ldots,z\}\quad(62\ \text{symbols}),$
+we exclude four visually ambiguous characters: lowercase $\ell$ and uppercase $\{I,J,O\}$. Thus,
+$|\mathcal{C}|=62-4=58.$
+0.B.1 Tasks and exact label spaces
+Numeric (two digits).
+The numeric label space contains all two-digit numbers without a leading zero:
+$\mathcal{N}=\{10,11,\ldots,99\},\qquad|\mathcal{N}|=90.$
+A uniform random-guess baseline is therefore
+$\Pr[\text{random correct}]=\frac{1}{|\mathcal{N}|}=\frac{1}{90}\approx 1.11\%.$
+Alphanumeric (two characters).
+The alphanumeric label space contains all ordered pairs from $\mathcal{C}$:
+$\mathcal{A}=\mathcal{C}\times\mathcal{C},\qquad|\mathcal{A}|=58^{2}=3364.$
+A uniform random-guess baseline is therefore
+$\Pr[\text{random correct}]=\frac{1}{|\mathcal{A}|}=\frac{1}{3364}\approx 0.0297\%.$
+0.B.2 Scoring policy
+We compute case-sensitive exact-match accuracy over $\mathcal{N}$ or $\mathcal{A}$, as applicable.
+Normalization prior to scoring trims whitespace, removes a possible “Answer: " prefix, extracts the first two alphanumerics.
+Characters outside $\mathcal{C}$ are considered invalid.
+Appendix 0.C A Brief Discussion on CIEDE2000 Color Difference
 $\Delta E_{2000}$ $=\sqrt{\left(\frac{\Delta L^{\prime}}{K_{L}S_{L}}\right)^{2}+\left(\frac{\Delta C^{\prime}}{K_{C}S_{C}}\right)^{2}+\left(\frac{\Delta H^{\prime}}{K_{H}S_{H}}\right)^{2}+R_{T}\left(\frac{\Delta C^{\prime}}{K_{C}S_{C}}\right)\left(\frac{\Delta H^{\prime}}{K_{H}S_{H}}\right)}$
-$\text{where}\quad&\Delta L^{\prime}\text{ is the corrected lightness difference,}\\ &\Delta C^{\prime}\text{ is the corrected chroma difference,}\\ &\Delta H^{\prime}\text{ is the corrected hue difference,}\\ &K_{L},K_{C},K_{H}\text{ are parametric factors (typically 1),}\\ &S_{L},S_{C},S_{H}\text{ are weighting functions for lightness, chroma, and hue,}\\ &R_{T}\text{ is a rotation term accounting for hue-chroma interaction.}\end{aligned}$ (1)
-The CIEDE2000 score ($\Delta E_{2000}$, Equation A) Luo et al.
-(2001) quantifies the perceived difference between two colors more accurately than prior formulae, especially for subtle variations.
+[…]
+The CIEDE2000 score ($\Delta E_{2000}$, Equation 0.C) [31] quantifies the perceived difference between two colors more accurately than prior formulae, especially for subtle variations.
 It calculates a single value representing the “distance” between colors in the perceptually uniform CIE $L^{*}a^{*}b^{*}$ space, considering lightness, chroma, and hue.
@@
 Figure 3: Distribution of CIEDE2000 color difference scores for the 25 selected foreground-background color pairs utilized in the HueManity benchmark.
-Appendix B Color Pairs Selection
+Appendix 0.D Color Pairs Selection
 The selection of appropriate color pairs for the foreground (characters) and background dots was a critical phase in the development of HueManity, undertaken with considerable care to ensure a balance between perceptual challenge and unambiguous human legibility.
@@
 This initial pool was iteratively refined by evaluating pairs against CIEDE2000 ($\Delta E_{2000}$, Eq.
-A) scores and visual checks.
+0.C) scores and visual checks.
 We retained promising candidates, modified some, and discarded others, while simultaneously manually crafting and vetting new pairs to meet the benchmark’s final requirements (detailed below).
@@
   2. 2.
-    Quantitative Contrast Filtering (CIEDE2000): Each of these candidate pairs then underwent rigorous quantitative analysis using the CIEDE2000 ($\Delta E_{2000}$) color difference formula (Equation A).
+    Quantitative Contrast Filtering (CIEDE 2000): Each of these candidate pairs then underwent rigorous quantitative analysis using the CIEDE2000 ($\Delta E_{2000}$) color difference formula (Equation 0.C).
 This formula is a standard measure in color science, designed to reflect perceptually meaningful differences as perceived by humans.
@@
 Therefore, we deliberately curated the final set of 25 color pairs to achieve an approximately equal distribution around a $\Delta E_{2000}$ score of 50.
-This threshold is grounded in color science literature, often considered a point distinguishing more subtle (scores <50) from more clearly distinct (scores >50) color differences.
+This threshold is grounded in color science literature, often considered a point distinguishing more subtle (scores << 50) from more clearly distinct (scores >> 50) color differences.
 We aimed for roughly half the selected pairs to fall below this threshold and half above, ensuring HueManity evaluates performance across varying, literature-informed degrees of color contrast difficulty.
   4. 4.
-    Manual Verification and Legibility Check: Recognizing that a single numerical contrast score like $\Delta E_{2000}$ captures overall perceived difference but may not fully account for the complex interplay of hue, saturation, and luminance components, especially when rendered as dots and subjected to further transformations (gradient, color, and light shifts as described in Section 3.2), a crucial final step of manual verification was performed.
-As you noted, it is hard to quantify the nuanced visual impact of these combined factors with a single metric.
-Therefore, for every color pair that passed the quantitative filtering, sample HueManity images were generated.
+    Manual Verification and Legibility Check: Recognizing that a single numerical contrast score like $\Delta E_{2000}$ captures overall perceived difference but may not fully account for the complex interplay of hue, saturation, and luminance components, especially when rendered as dots and subjected to further transformations (gradient, color, and light shifts), a crucial final step of manual verification was performed.
+It is hard to quantify the nuanced visual impact of these combined factors with a single metric, therefore for every color pair that passed the quantitative filtering, sample HueManity images were generated.
 These renderings were meticulously inspected by the authors.
@@
 This ensures that the stimuli used in HueManity are not only theoretically sound but also practically validated for fairness, legibility, and the intended level of perceptual challenge.
-Appendix C Qualitative Analysis of MLLM Failure Patterns
+Table 7: The 25 Curated Color Pairs Used in the HueManity Benchmark.
+Foreground Color Background Color (Foreground | Background) RGB Values
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+Appendix 0.E Qualitative Analysis of MLLM Failure Patterns
 This section details common failure patterns observed in Multimodal Large Language Models (MLLMs) when tasked with identifying alphanumeric characters embedded in Ishihara-style dot patterns from the HueManity dataset.
@@
 In contrast, MLLMs exhibited distinct and more fundamental failure modes.
-C.1 Prevalent Hallucination of Unrelated or Overly Complex Characters
+0.E.1 Prevalent Hallucination of Unrelated or Overly Complex Characters
 A dominant failure mode across multiple MLLMs was the generation of characters, words, or even entire phrases that bore no resemblance to the two-character ground truth.
@@
 This pattern suggests that when the fine-grained perceptual challenge overwhelms the MLLMs’ visual processing, they may default to generating text that, while perhaps linguistically plausible, is detached from the actual visual content.
-C.2 Frequent Resort to Descriptive Evasion or Explicit Admission of Inability
+0.E.2 Frequent Resort to Descriptive Evasion or Explicit Admission of Inability
 Rather than consistently attempting to identify the embedded characters, many MLLMs frequently defaulted to one of two evasive strategies: providing a general description of the image (often correctly identifying it as a color vision test) or explicitly stating their incapacity to discern any characters.
@@
 This pattern suggests that MLLMs may possess internal confidence thresholds that, when triggered by low-confidence visual parsing, lead to evasive or pre-programmed “unable to process” responses rather than a forced, best-guess attempt at character recognition.
-C.3 Erratic, Unpredictable, and Systematically Flawed Output Patterns
+0.E.3 Erratic, Unpredictable, and Systematically Flawed Output Patterns
 MLLM outputs were frequently characterized by their erratic and unpredictable nature.
@@
 This high variance in error types was observed both across different models for the same input and within the outputs of a single model across different images.
-For instance, when presented with the same alphanumeric target (e.g., “Wh”), while one model (GPT-4.1) might respond almost correctly, others exhibited diverse failures: Claude 3.7 Sonnet produced an unrelated number (“4726”); LLaVA-13b generated an exceptionally long string of sequential numbers; and Qwen VL Max incorrectly reasoned the presence of a different number (“12”).
+For instance, when presented with the same alphanumeric target (e.g., “Wh”), while one model (GPT-4.1) might respond almost correctly, others exhibited diverse failures: Claude 3.7 Sonnet produced an unrelated number (“4726”), LLaVA-13b generated an exceptionally long string of sequential numbers, and Qwen VL Max incorrectly reasoned the presence of a different number (“12”).
 Some incorrect outputs also suggested flawed systematic processing, such as LLaVA-13b responding with a patterned string like “[L1L1L1]” for one target or generating extremely long, patterned numeric strings for others.
@@
 This unpredictability underscores a lack of robust and stable visual feature extraction and interpretation, contrasting with human visual processing, which tends towards predictable errors based on similarity.
+Appendix 0.F Usage of Generative AI tools
+We utilized Generative AI tools to help improve the language, phrasing, and readability of this manuscript.